OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? a + 98 = -2

OpenStudy (anonymous):

add -98 on both sides

OpenStudy (anonymous):

ok bt is the answer 96 but i have to slove it can u do it for me plz

OpenStudy (anonymous):

a+98=-2 Adding (-98) on both sides, a+98+ (-98) = -2 + (-98) a+98 -98 = -2-98 a=-100

OpenStudy (anonymous):

but what i didnt do get it

OpenStudy (anonymous):

but that was correct

OpenStudy (anonymous):

what answer did you get ?

OpenStudy (anonymous):

same as u

OpenStudy (anonymous):

so now you understand ?

OpenStudy (anonymous):

kinda og

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

do you have any doubts ?

OpenStudy (anonymous):

what